OpenStudy (anonymous):

HELP ME PLEASE"!! how do you solve this step by step sqr(x+1)=sqr(x+1)

OpenStudy (anonymous):

its all real numbers. its an identity

OpenStudy (anonymous):

but the anwer the books gives me is 0

OpenStudy (anonymous):

as in (X+1)^2=(X+1)^2?

OpenStudy (anonymous):

no is square root (x+1)= square root (x+1) it says solving radical equations

OpenStudy (anonymous):

yah still just means a=a; it is an identity which is always true

OpenStudy (anonymous):

is 0 the same thing as all real numbers then?

OpenStudy (anonymous):

no, the book is wrong, 0 is just one of an infinite set of solutions

OpenStudy (anonymous):

oh ok thank you :)

OpenStudy (anonymous):

oh ok thank you :)